# MANGO: A Benchmark for Evaluating Mapping and Navigation Abilities of Large Language Models

### v1

@@ 2.2 Generation of Question Skeletons: Traversing Mazes and Imputing Edges
After extending all the mazes in our benchmark, we collected 21046 DF skeletons and 14698 RF skeletons by traversing the extended graphs. Being evaluated on a maze, the LLM may not be able to consume the entire walkthrough in its context window. That is, we may only feed it an appropriate prefix of the walkthrough (e.g., the first $70$ steps for Zork-I as shown in LABEL:lst:walkthrough), leaving some of the DF and RF skeletons unanswerable given that prefix. Therefore, our benchmark provides the ANSWERABLE label (an integer) for each skeleton such that this skeleton is only answerable if the maximum STEP NUM in that prefix (e.g., $70$ in LABEL:lst:walkthrough) is greater than or equal to its ANSWERABLE label. Furthermore, given a walkthrough prefix, an answerable skeleton may be easy or hard, depending on whether it involves edges that are not covered in the prefix. Precisely, a DF skeleton is considered to be easy if all the S-A-D edges in its corresponding simple path are covered in the walkthrough prefix; an RF skeleton is easy if the shortest simple path from its starting location to its destination only involves the S-A-D steps covered in the prefix. When a longer walkthrough prefix is used, more answerable questions tend to become easy. Our benchmark provides the EASY label (also an integer) for each skeleton: a skeleton is easy if the maximum STEP NUM in the walkthrough prefix is no smaller than its EASY label; otherwise, it is a hard skeleton. Table 2 in Appendix A documents the statistics of the full dataset, such as the number of locations and the number of skeletons. Tables 4, 7, 5 and 6 in Appendix B shows the information about the data on which each LLM was evaluated in our experiments.
@@ 3.2 Main Results
More results are in Appendix C, including results on other evaluation metrics (e.g., weighted success rates) and comparison between Llama-2 with Llama-1 and Llama-2-chat.
@@ 3.4 Does Mapping and Navigation Ability Matter in Downstream Tasks?
Now we present a case study showing that a strong mapping and navigation ability of an LLM would benefit it in downstream tasks. In particular, we selected 284 minigames in the Jericho game suite, and investigated how the map knowledge may improve the performance of an LLM in playing these minigames. Each minigame is a selected step from one of 53 textgames; the selection criterion is that the best action to take at this step is a movement. Through evaluating an LLM on a minigame, we synthesize a scenario in which the LLM has to figure out the best action to take given its previous actions and observations (i.e., the prefix of walkthrough up to the current step). Note that this task is different and more challenging than answering the DF and RF questions: the LLM is not explicitly given a route (as in DF questions) or a destination (as in RF questions), but has to spontaneously figure out which action may contribute to its long-term goal.
@@ 4 Related Work
Over the past few years, the field of natural language processing has experienced remarkable advancements with the emergence of large language models. This progress has spurred a multitude of research endeavors that propose benchmarks challenging the limits of these models. Those benchmarks assess the capacities of LLMs in linguistics (Wang et al., 2018, 2019), reading comprehension (Richardson et al., 2013; Lai et al., 2017), commonsense reasoning (Zellers et al., 2019; Bisk et al., 2020; Huang et al., 2019; Talmor et al., 2019), arithmetic reasoning (Miao et al., 2020; Cobbe et al., 2021; Patel et al., 2021), and knowledge memorization and understanding (Clark et al., 2018; Mihaylov et al., 2018; Khot et al., 2020; Clark et al., 2020; Hendrycks et al., 2021; Srivastava et al., 2022). Recent models have achieved remarkable performance not only on these benchmarks, but also across a diversity of human-oriented academic and professional exams (OpenAI, 2023) as well as general tasks (Bubeck et al., 2023). Our benchmark presents a unique challenge to large language models, evaluating their capacity to acquire spatial knowledge about new environments and answering complex navigation questions; it is a dimension orthogonal to the aforementioned reasoning abilities.
@@ Appendix A Benchmark Details
As introduced in section 2.2, each skeleton is paired with an ANSWERABLE label and an EASY label: given a prefix of the walkthrough, the ANSWERABLE label indicates whether this skeleton is answerable, and the EASY label will decide whether this skeleton is easy or hard given this walkthrough prefix. Table 2 displays some important statistics of the full dataset broken down into each individual maze, including the number of DF and RF skeletons. In addition, we also provide the following data files for easy reference:
@@
905 5 7 1.88 21 11 5 11 5
advent 70 137 19.38 276 24788 4220 4694 0
adventureland 23 48 7.46 169 1117 208 484 22
afflicted 13 24 3.18 97 156 0 156 0
anchor 118 228 27.46 530 66133 23564 13573 0
awaken 15 28 5.02 56 365 45 171 25
balances 12 20 3.22 121 136 1 94 1
ballyhoo 39 90 9.26 415 18148 6298 1444 0
cutthroat 52 108 10.76 335 5969 2688 1832 59
deephome 66 127 8.64 326 4161 235 4161 1
detective 32 41 8.76 50 505 6 505 6
dragon 24 52 8.9 100 1100 2262 529 23
enchanter 57 110 10.93 264 5118 2602 2773 99
enter 18 34 3.88 101 306 0 306 0
gold 22 44 4.99 344 682 0 462 0
hhgg 42 65 7.02 360 3794 3 1602 1
hollywood 109 219 27.33 396 21641 15557 11451 213
huntdark 12 13 4.24 66 66 2 66 2
infidel 58 122 9.27 249 2786 1729 2158 453
inhumane 43 91 6.49 121 1623 3020 1474 332
jewel 43 74 8.66 222 1157 61 1157 61
karn 56 124 15.36 361 24479 108983 3025 0
library 7 12 2.48 51 42 0 42 0
loose 12 21 4.18 49 94 27 94 27
lostpig 13 26 4.5 145 492 105 114 0
ludicorp 86 176 14.96 363 7099 7281 6002 1308
lurking 60 116 13.04 293 5187 2316 2946 105
moonlit 6 9 2.2 58 18 7 18 7
murdac 84 157 11.05 303 6914 2016 5967 1005
night 20 41 6.93 89 633 59 380 0
omniquest 32 65 7.55 77 648 1431 642 90
partyfoul 4 8 1.92 55 24 0 12 0
pentari 18 31 3.76 48 208 20 208 4
planetfall 69 138 12.41 398 7100 2246 3887 247
plundered 45 87 13.46 188 3705 2394 1393 26
reverb 17 31 5.26 73 321 20 256 16
seastalker 20 43 7.55 203 1123 1143 260 3
sherlock 71 140 12.39 338 16417 11758 4214 27
snacktime 4 6 1.5 33 12 0 12 0
sorcerer 64 120 11.72 253 4042 2285 2503 264
spellbrkr 73 111 13.62 411 8713 1762 3462 1
spirit 229 466 26.39 1263 111102 47410 51531 226
temple 24 46 6.07 180 563 106 486 21
theatre 20 38 5.33 295 329 130 329 51
trinity 94 189 20.39 609 65049 131324 8742 0
tryst205 73 139 10.04 517 5089 392 3983 181
wishbringer 46 94 13.32 183 4988 5646 2070 0
yomomma 9 25 3.55 97 189 263 64 0
zenon 17 32 4.85 82 144 128 144 128
zork1 84 166 20.79 395 20984 26857 6889 0
zork2 66 136 11.79 295 15174 13081 3798 178
zork3 55 105 15.37 272 6202 2754 2294 133
ztuu 16 30 3.27 83 225 15 225 15
Table 2: Statistics of full data. Here, # Locs represents the number of locations in each walkthrough; # Edges represents the number of edges; AVG LEN PATH denotes the average length of all paths; # STEPS indicates the number of steps in each walkthrough; EASY and HARD of the DF and RF respectively represent the number of easy and hard skeletons of the DF and RF tasks. When counting the easy and hard skeletons, we assume that the full walkthrough will be used. For the statistics of the data used in our experiments, please see Tables 4, 7, 6 and 5.
@@ A.2 Location Resolution Details
  • add the machine ID (e.g., ???) to the location name as a mark (e.g., ??? to ???);
  • if the location name already has a mark (e.g., 1st floor, north end of north/south hall, thanks to the name resolution process), we replace the mark with the machine ID (e.g., ???).
@@ B.2 Prompt Details
What prompts are to LLMs are like what hyperparameters are to classifical deep neural nets. To ensure that we could obtain the optimal results of RWKV and Llamas, we tuned the prompts—more precisely, experimented with variants of the prompts of GPTs—and ended up with a set of new prompts that are mostly the same but exhibit some prose differences. For example, at the beginning of the prompts, we added “Here is a walkthrough of a text game.”
It is worth noting that, even though we carefully tuned the prompts, the non-GPT models still suffer a high chance of failing to return well-structured answers. Table 3 shows the average number of answered questions as well as how many of them received ill-formatted answers. As we could see, GPT-3.5 and GPT-4 could generate well-structured answers for a large portion of the DF and RF questions, but the other models often gave ill-structured answers. As a result, we could only evaluate the non-GPT models on the questions to which the answers were well-structured and thus could be parsed. We also experimented with the function-calling interface of GPTs, but it didn’t lead to an increased amount of well-structured answers compared to our prompt design.
@@ B.3 Walkthrough and Question Details
Due to the context window size of an LLM, it is often the case that we have to only use a prefix of the walkthrough when evaluating an LLM on a maze. Recall from section 2.2 that not every question is answerable given a walkthrough prefix. Therefore, each model was evaluated on a different set of questions. Tables 4, 7, 6 and 5 display the statistics about the data that each LLM was actually evaluated on in our experiments.
@@ C.4 Results of Search-Based Approaches
In both settings, GPT-3.5 yields a low accuracy of the S-A-D triplet completion, indicating the difficulty of translating natural language into symbolic graphs. In the first setting, the average accuracy is 70.5 %. In the second setting, the average success rate is 6.3 %. The second setting is more challenging because any mistake would cause all subsequent steps to be incorrect. Some LLM mistakes and our error analysis can be found at https://hackmd.io/@anonymous123123123/BJJvjGhVT
@@ C.5 More Llama Results
C.6 More Results About Playing Minigames
@@ Appendix D Future Directions
Our MANGO benchmark sets up a leaderboard, measuring the mapping and navigation abilities of LLMs. An interesting future direction is to investigate how the internal representations of each LLM have—if any—captured the structures of the maps. This can potentially reveal critical insights into the underlying mechanism of an LLM to represent spatial relations, benefiting their application into real-world scenarios.
@@
Such upgrades will enhance its practical application in more complex scenarios.

### v2

@@ 2 MANGO: A Benchmark for Text-Based Mapping and Navigation
author=DP,color=blue!20,size=,fancyline,caption=,]in rebuttal, we promised a complete example of walkthrough, we may use partial zork1 as appendix or some other ones with less step
@@ 2.2 Generation of Question Skeletons: Traversing Mazes and Imputing Edges
After extending all the mazes in our benchmark, we collected 21046 DF skeletons and 14698 RF skeletons by traversing the extended graphs. Being evaluated on a maze, the LLM may not be able to consume the entire walkthrough in its context window. That is, we may only feed it an appropriate prefix of the walkthrough (e.g., the first $70$ steps for Zork-I as shown in LABEL:lst:walkthrough), leaving some of the DF and RF skeletons unanswerable given that prefix. Therefore, our benchmark provides the ANSWERABLE label (an integer) for each skeleton such that this skeleton is only answerable if the maximum STEP NUM in that prefix (e.g., $70$ in LABEL:lst:walkthrough) is greater than or equal to its ANSWERABLE label. Furthermore, given a walkthrough prefix, an answerable skeleton may be easy or hard, depending on whether it involves edges that are not covered in the prefix. Precisely, a DF skeleton is considered to be easy if all the S-A-D edges in its corresponding simple path are covered in the walkthrough prefix; an RF skeleton is easy if the shortest simple path from its starting location to its destination only involves the S-A-D steps covered in the prefix. When a longer walkthrough prefix is used, more answerable questions tend to become easy. Our benchmark provides the EASY label (also an integer) for each skeleton: a skeleton is easy if the maximum STEP NUM in the walkthrough prefix is no smaller than its EASY label; otherwise, it is a hard skeleton. Table 3 in Appendix A documents the statistics of the full dataset, such as the number of locations and the number of skeletons. Tables 5, 8, 6 and 7 in Appendix B shows the information about the data on which each LLM was evaluated in our experiments.
@@ 3.2 Main Results
More results are in Appendix C, including results on other evaluation metrics (e.g., weighted success rates) and comparison between Llama-2 with Llama-1 and Llama-2-chat. We also explored an alternative approach that first maps a walkthrough to a symbolic graph and then uses a search algorithm to answer the given (DF or RF) question. Results and analysis of this approach can be found in section C.4. Overall, we found it to be very challenging to translate natural language walkthroughs into searchable symbolic graphs in the first place, making this approach not promising.
@@ 3.3 Analysis of GPTs
3.4 Human Performance
We measured human performance on a subset of our data. This subset includes 30 DF questions (21 easy and 9 hard) and 31 RF questions (20 easy, 11 hard). The student authors participated in two rounds of evaluation: in the first round, each author answered a random split of the questions, and all the questions were answered; in the second round, we randomly sampled 10 DF and 10 RF questions and let each of them be re-evaluated by a different author. The second round allows us to analyze human agreement; this analysis can be found in section C.6.
Task Difficulty Success Rate Reasoning Accuracy
Route Finding All (31) 0.8211 0.6129
Easy (20) 0.7727 0.5500
Hard (11) 0.9091 0.7273
Destination Finding All (30) 1.0000 0.5667
Easy (21) 1.0000 0.6667
Hard (9) 1.0000 0.3333
Table 2: Human Evaluation Results
As shown in Table 2, humans generally achieve high success rates on both RF and DF questions, and exhibit small to no difference across easy and hard questions. This is interesting since LLMs tend to struggle on hard questions. Reasoning accuracies are lower than success rates; this is because human answers exhibit a relatively high variability in writing (e.g., location names with special symbols).
Despite the high scores, answering the questions is actually not trivial for human raters. Answering each DF question takes an average of 15 minutes, while answering each RF question takes an average of 30 minutes. All the human raters needed to take notes (e.g., drawing maps) while completing these tasks.
@@ 3.5 Does Mapping and Navigation Ability Matter in Downstream Tasks?
Now we present a case study showing that a strong mapping and navigation ability of an LLM would benefit it in downstream tasks. In particular, we selected 284 minigames in the Jericho game suite, and investigated how the map knowledge may improve the performance of an LLM in playing these minigames. Each minigame is a selected prefix of a walkthrough from one of 53 textgames; the selection criterion is that the best action to take at this step is a movement. In other words, each minigame is a scenario where the LLM has to figure out the best action to take given its previous actions and observations (i.e., the prefix of walkthrough up to the current step). This task is different and more challenging than answering the DF and RF questions: the LLM is not explicitly given a route (as in DF questions) or a destination (as in RF questions), but has to spontaneously figure out which action may contribute to its long-term goal.
@@ 4 Related Work
Over the past few years, the field of natural language processing has experienced remarkable advancements with the emergence of large language models. This progress has spurred a multitude of research endeavors that propose benchmarks challenging the limits of these models. Those benchmarks assess the capacities of LLMs in linguistics (Wang et al., 2018; 2019), reading comprehension (Richardson et al., 2013; Lai et al., 2017), commonsense reasoning (Zellers et al., 2019; Bisk et al., 2020; Huang et al., 2019; Talmor et al., 2019), arithmetic reasoning (Miao et al., 2020; Cobbe et al., 2021; Patel et al., 2021), and knowledge memorization and understanding (Clark et al., 2018; Mihaylov et al., 2018; Khot et al., 2020; Clark et al., 2020; Hendrycks et al., 2021; Srivastava et al., 2022). Recent models have achieved remarkable performance not only on these benchmarks, but also across a diversity of human-oriented academic and professional exams (OpenAI, 2023) as well as general tasks (Bubeck et al., 2023). Our benchmark presents a unique challenge to large language models, evaluating their capacity to acquire spatial knowledge about new environments and answering complex navigation questions; it is a dimension orthogonal to the aforementioned reasoning abilities.
@@ 5 Conclusion
Acknowledgments
This work was supported by a research gift to the last author by Adobe Research. It was also supported by the TRI University 2.0 Project, the U.S. National Science Foundation (Grant Number CNS-1948457 and CNS-2340171), and Office of Advanced Scientific Computing Research at Department of Energy (Award Number DE-SC0024424). The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof. We thank the anonymous COLM reviewers and meta-reviewer for their constructive feedback. We also thank our colleagues at TTIC and UChicago for helpful discussion.
@@ Appendix A Benchmark Details
As introduced in section 2.2, each skeleton is paired with an ANSWERABLE label and an EASY label: given a prefix of the walkthrough, the ANSWERABLE label indicates whether this skeleton is answerable, and the EASY label will decide whether this skeleton is easy or hard given this walkthrough prefix. Table 3 displays some important statistics of the full dataset broken down into each individual maze, including the number of DF and RF skeletons. In total, our full dataset has about 3M DF questions and 200K RF questions. On average, each maze has around 60K DF questions and 4K RF questions. Noticeably, the 53 games in our dataset are very diverse: they cover a range of topics and genres; they are situated in different eras; they cover a wide variety of maps (small vs. big houses, long vs. short halls, towns vs. forests, verbose vs. concise scene descriptions, etc). In addition, we also provide the following data files for easy reference:
@@
905 5 7 1.88 22 11 5 11 5
advent 70 137 19.38 277 24788 4220 4694 0
adventureland 23 48 7.46 170 1117 208 484 22
afflicted 13 24 3.18 99 156 0 156 0
anchor 118 228 27.46 531 66133 23564 13573 0
awaken 15 28 5.02 57 365 45 171 25
balances 12 20 3.22 122 136 1 94 1
ballyhoo 39 90 9.26 416 18148 6298 1444 0
curses 150 304 33.83 816 1951931 549433 22201 0
cutthroat 52 108 10.76 336 5969 2688 1832 59
deephome 66 127 8.64 327 4161 235 4161 1
detective 32 41 8.76 51 505 6 505 6
dragon 24 52 8.9 101 1100 2262 529 23
enchanter 57 110 10.93 265 5118 2602 2773 99
enter 18 34 3.88 102 306 0 306 0
gold 22 44 4.99 345 682 0 462 0
hhgg 42 65 7.02 361 3794 3 1602 1
hollywood 109 219 27.33 397 21641 15557 11451 213
huntdark 12 13 4.24 67 66 2 66 2
infidel 58 122 9.27 250 2786 1729 2158 453
inhumane 43 91 6.49 122 1623 3020 1474 332
jewel 43 74 8.66 223 1157 61 1157 61
karn 56 124 15.36 362 24479 108983 3025 0
library 7 12 2.48 52 42 0 42 0
loose 12 21 4.18 50 94 27 94 27
lostpig 13 26 4.5 146 492 105 114 0
ludicorp 86 176 14.96 364 7099 7281 6002 1308
lurking 60 116 13.04 294 5187 2316 2946 105
moonlit 6 9 2.2 59 18 7 18 7
murdac 84 157 11.05 304 6914 2016 5967 1005
night 20 41 6.93 90 633 59 380 0
omniquest 32 65 7.55 78 648 1431 642 90
partyfoul 4 8 1.92 56 24 0 12 0
pentari 18 31 3.76 49 208 20 208 4
planetfall 69 138 12.41 399 7100 2246 3887 247
plundered 45 87 13.46 189 3705 2394 1393 26
reverb 17 31 5.26 74 321 20 256 16
seastalker 20 43 7.55 204 1123 1143 260 3
sherlock 71 140 12.39 339 16417 11758 4214 27
snacktime 4 6 1.5 34 12 0 12 0
sorcerer 64 120 11.72 254 4042 2285 2503 264
spellbrkr 73 111 13.62 412 8713 1762 3462 1
spirit 229 466 26.39 1264 111102 47410 51531 226
temple 24 46 6.07 181 563 106 486 21
theatre 20 38 5.33 296 329 130 329 51
trinity 94 189 20.39 610 65049 131324 8742 0
tryst205 73 139 10.04 518 5089 392 3983 181
wishbringer 46 94 13.32 184 4988 5646 2070 0
yomomma 9 25 3.55 98 189 263 64 0
zenon 17 32 4.85 83 144 128 144 128
zork1 84 166 20.79 396 20984 26857 6889 0
zork2 66 136 11.79 296 15174 13081 3798 178
zork3 55 105 15.37 273 6202 2754 2294 133
ztuu 16 30 3.27 84 225 15 225 15
Table 3: Statistics of full data. Here, # Locs represents the number of locations in each walkthrough; # Edges represents the number of edges; AVG LEN PATH denotes the average length of all paths; # STEPS indicates the number of steps in each walkthrough; EASY and HARD of the DF and RF respectively represent the number of easy and hard skeletons of the DF and RF tasks. When counting the easy and hard skeletons, we assume that the full walkthrough will be used. For the statistics of the data used in our experiments, please see Tables 5, 8, 7 and 6.
@@ A.2 Location Resolution Details
  • add the machine ID (e.g., “(obj59)”) to the location name as a mark (e.g., “in debris room” to “in debris room (obj59)”);
  • if the location name already has a mark (e.g., 1st floor, north end of north/south hall, thanks to the name resolution process), we replace the mark with the machine ID (e.g., “hall (1st floor, middle of north/south hall)” to “hall (obj66)”).
@@ B.2 Prompt Details
What prompts are to LLMs are like what hyperparameters are to classifical deep neural nets. Our experiments show that the structure of LLM output is sensitive to the prompts, and thus we had to carefully tune the prompts such that LLMs could return easy-to-parse output. For example, we found it helpful to ask LLMs to format their answers as a Python list of Python dictionaries; we also found it helpful to end our prompt with the symbol ‘[’—i.e., the start symbol of a Python list—to elicit the LLM to actually output in the desired format. Fortunately, as long as the output follows the desired structure, its content is relatively robust to the prompts. In our pilot experiments, we found that different prompts (created by different authors) would generate the same or similar content (regardless of its format) and achieve the same level of success rates (after parsing).
To ensure that we could obtain the optimal results of RWKV and Llamas, we tuned the prompts—more precisely, experimented with variants of the prompts of GPTs—and ended up with a set of new prompts that are mostly the same but exhibit some prose differences. For example, at the beginning of the prompts, we added “Here is a walkthrough of a text game.” It is worth noting that, even though we carefully tuned the prompts, the non-GPT models still suffer a high chance of failing to return well-structured answers. Table 4 shows the average number of answered questions as well as how many of them received ill-formatted answers. As we could see, GPT-3.5 and GPT-4 could generate well-structured answers for a large portion of the DF and RF questions, but the other models often gave ill-structured answers. As a result, we could only evaluate the non-GPT models on the questions to which the answers were well-structured and thus could be parsed. We also experimented with the function-calling interface of GPTs, but it didn’t lead to an increased amount of well-structured answers compared to our prompt design.
@@ B.3 Walkthrough and Question Details
Due to the context window size of an LLM, it is often the case that we have to only use a prefix of the walkthrough when evaluating an LLM on a maze. Recall from section 2.2 that not every question is answerable given a walkthrough prefix. Therefore, each model was evaluated on a different set of questions. Tables 5, 8, 7 and 6 display the statistics about the data that each LLM was actually evaluated on in our experiments.
@@ C.4 Results of Search-Based Approaches
In both settings, GPT-3.5 yields a low accuracy of the S-A-D triplet completion, indicating the difficulty of translating natural language into symbolic graphs. In the first setting, the average accuracy is 70.5 %. In the second setting, the average success rate is 6.3 %. The second setting is more challenging because any mistake would cause all subsequent steps to be incorrect. Some LLM mistakes and our error analysis can be found at https://github.com/Oaklight/mango/tree/camera-ready/utils/supp_exp.
@@ C.5 More Llama Results
C.6 Human Performance Details
To measure human agreement, we computed the mean square error (MSE) for each metric across the questions that were answered by more than one human rater. As shown in Table 13, the human agreement turns out to be high.
All Easy Hard
Task Type SR RA SR RA SR RA
Route Finding 0.1442 0.3333 0.0425 0.1429 0.5000 1.0000
Destination Finding 0.0000 0.6000 0.0000 0.5714 0.0000 0.6667
Table 13: Human agreement measured by MSE. SR stands for success rate, while RA denotes reasoning accuracy.
C.7 More Results About Playing Minigames
@@ Appendix D Future Directions
Our MANGO benchmark sets up a leaderboard, measuring the mapping and navigation abilities of LLMs. An interesting future direction is to investigate how the internal representations of each LLM have—if any—captured the structures of the maps. For example, one may probe the LLM representations and examine whether these representations are predictive of certain characteristics of the maze, similar to the experiments of Li et al. (2023) in board games. This can potentially reveal critical insights into the underlying mechanism of an LLM to represent spatial relations, benefiting their application into real-world scenarios.
@@
Such upgrades will enhance its practical application in more complex scenarios. They will also enable investigation in interesting topics, including
  • LLMs’ generalization to significantly out-of-distribution navigation scenarios (e.g., different structures, styles, and configurations);
  • LLMs’ robustness to distracting context.